OpenStudy (anonymous):

If cos θ = 4/5, what is cos (- θ)?

OpenStudy (anonymous):

\[\cos (- \theta)=\cos(\theta)= 4/5\]\[\theta \]

OpenStudy (anonymous):

yeah i know you switch it, but now do I just plug that in my calculator or leave it?

OpenStudy (anonymous):

just for 4/5

OpenStudy (anonymous):

cosine is even function, f(-x) = f(x)

OpenStudy (anonymous):

ohkay so it would be cos(- θ)= 4/5?

OpenStudy (anonymous):

Try drawing the two triangles on a coordinate grid. 

OpenStudy (anonymous):

ohkay, im not sure how that helps me?

OpenStudy (anonymous):

Do you know what the definition of cosine is?

OpenStudy (anonymous):

no

OpenStudy (aravindg):

cos(-theta)=cos theta

OpenStudy (anonymous):

so then its just 4/5

OpenStudy (aravindg):

yep!!!

OpenStudy (anonymous):

thanks!!